A 55.33 g KNO3 with molar mass of 101 g/mole was placed in 141g of H2O. with molar mass of 18.0 g/mole. What is the percent by mass of solution?


1
Expert's answer
2021-03-17T06:39:51-0400

Molar mass of KNO3 = 101 g/mol

Mass of KNO3 = 55.33 g

Moles of KNO3"(n_1)=" "\dfrac{55.33}{101}=0.547\space mol"

Molar mass of H2O = 18 g/mol

Mass of H2O = 141 g

Moles of H2O"(n_2)=\dfrac{18}{141}=0.127\space mol"

Mole fraction of KNO3 "=\dfrac{n_1}{n_1+n_2}=0.81"

Mole fraction of H2O "=0.19"

Total mass of solution = 196.33 g

Average Molecular mass of solution"(M_{av})=\sum M_i x_i=0.81\times 101+0.19\times18"

"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=85.23 \space g"

Percent by weight = "\dfrac{85.23}{196.33}\times 100=43.41\%"